Show algebraically that the circles whose equations are x^2+y^2=16 and
x^2+y^2-20x+64=0 are externally tangent. Find the area of their point of tangency.

Answer by Alan3354(69443)

Show algebraically that the circles whose equations are x^2+y^2=16 and
x^2+y^2-20x+64=0 are externally tangent. Find the area of their point of tangency.
----------
x^2+y^2=16
This circle is about the Origin and r = 4
----------------
x^2+y^2-20x+64=0
Complete the square:
x^2-20x+100 + y^2+64=100
(x-10)^2 + y^2 = 36
This circle's center is (10,0) and r = 6
---------------
The 2 radii add to 10, the same as the distance between the centers so they're tangent.
The point of tangency is (4,0). They're tangent, there's no area involved.